Assuming no technological progress but population and efficiency of labor growth per period at rates, n and g respectively. Given the Cobb-Douglas production function:

Y equals K to the power of alpha open parentheses E L close parentheses to the power of 1 minus alpha end exponent

a. Characterize the steady-state capital per effective worker

b. Characterize the steady-state output per effective worker

c. Determine a. and b. when a = .3

; delta space equals space 0.03 ; n equals 0.02 ;  g equals 0.01  and s equals 0.24 .
Answer #1

@ Y = X 1 ** (LE)!-( = K (LE) * = kx (LEY CL LE be E) In steady state sy- (stn+g) k =0 sy = (stn+g)R s(R2Q = (S+n+7) R.stntg Re stntg S 70 st ntg (6) y = stntg GC Stnrg x= 0.3 n = S = g: s= 0.02 0.03 0.01 0.24 stntg 1-0.3 0.24 0.03 +0.02 +0.01R= 0:29) 4. 251 = * 0.3 | نه 8: | | | مه |عے ل والية